# On the common index jump theorem and further developments

### v1

@@ 3 The coincidence of ${\beta}_{-}({\gamma})$ in [CGG24] and $S_{{\gamma}(1)}^{-}(1)$ in [Lo02]
Note that such identities were first introduced in [Lo90] of 1990 and rigorously proved in [Lo97] of 1997. Then they yield
@@
To prove (IR1”-1), letting $k$ and $d$ in (IR1”-1) be given by $k=2m_{i}$ and $d=2N$. Note first that using notations in [Lo02] and [DLW16], the left hand side of (IR1”-1) is given by
@@
where $2m_{i}\hat{i}(\Phi_{i})=\hat{i}(\Phi_{i}^{2m_{i}})$ holds. Then by (3.42) and (3.43) of [DLW16], for any given $\eta>0$, we obtain
@@
Next using notations in [Lo02] and [DLW16], the integer $d$ in (IR1”-2) is given by $d=2N$ too. Therefore (IR1”-2) coincides precisely with (1) of Theorem 1.1 of [LL00] (cf. $1^{o}$ of Theorem 10.1.2 on p.213 of [Lo02]).
@@
5 The coincidence of (IR2) and (IR3) in Theorem 3.3 of [CGG24] and (3.20)-(3.22) in Theorem 3.5 of [DLW16]
Now in the studies of the next two propositions, we prove that (IR2) and (IR3) of [CGG24] coincide completely with (3.20)-(3.22) of Theorem 3.5 in [DLW16]. In these proofs we use the notations in [DLW16] and [Lo02].
Proposition 3. The assertion (IR2”) and then (IR2) in [CGG24] coincides precisely with (3.20)-(3.21) in [DLW16].
Proof. Note first by Corollary 6.1.12 and Definition 6.1.13 on p.144 of [Lo02], it yields immediately (4.3)-(4.5). Therefore (IR2”) and then (IR2) in [CGG24] coincides precisely with (3.20)-(3.21) of [DLW16].
@@
Proposition 4. The assertion (IR3”) and then (IR3) in [CGG24] coincides precisely with (3.20)-(3.22) in [DLW16].
Proof. In order to use uniformly the notations in [DLW16] in the following proof, we have changed the notations ${\bf R}^{2m}$, ${\rm Sp}(2m)$, $\ell$ and $\Phi$ in (IR3”) of [CGG24] to ${\bf R}^{2n}$, ${\rm Sp}(2n)$, $m$ and $\gamma$} as those in (3.20) and (3.22) of [DLW16] respectively,
Note first that by the requirement of Theorem 3.3 of [CGG24] on the choice of $k$, the degree $D$ of ${\gamma}_{k}$ is a factor of $k$, and by the requirement (3.30) on $m_{k}$ of Theorem 3.5 of [DLW16], there holds $\frac{m_{k}{\theta}}{\pi}\in{\bf Z}$ whenever $\frac{{\theta}}{\pi}\in{\bf Q}\cap(0,2)$ and $e^{\sqrt{-1}{\theta}}\in{\sigma}(M_{k})$. These two requirements are precisely the same.
By (IR3”) in [CGG24] and (3.22) in [DLW16], to prove Proposition 4, it suffices to prove the following identity for every symplectic path in $C([0,1],{\rm Sp}(2n))$,
@@
where $M_{k}={\gamma}_{k}(1)$, and $Q_{k}(m)$ is given by [DLW16,(3.25)].
@@
where the last equality follows from [DLW16, (3.20)].
@@
where to get the last equality we have applied the fact $S_{M_{k}}^{+}(1)=S_{M_{k}}^{-}(1)$ from Lemma 9.1.6 on p.192 of [Lo02].
Then by the Bott-type formula for splitting numbers which was first proved in [LLZ02] (cf. Corollary 9.2.4 on p.201 in [Lo02]), we obtain

### v2

@@ 1 Introduction
Note that such identities were first introduced in [10] of 1990 and rigorously proved in [11] of 1997 by Long. Then they yield
@@
To prove (IR1”-1), letting $k$ and $d$ in (IR1”-1) be given by $k=2m_{i}$ and $d=2N$. Note first that using notations in [15] and [3], the left hand side of (IR1”-1) is given by
@@
where $2m_{i}\hat{i}(\Phi_{i})=\hat{i}(\Phi_{i}^{2m_{i}})$ holds. Then by (3.42) and (3.43) of [3], for any given $\eta>0$, we obtain
@@
Next using notations in [15] and [3], the integer $d$ in (IR1”-2) is given by $d=2N$ too. Therefore (IR1”-2) coincides precisely with (1) of Theorem 1.1 of [8] (cf. $1^{0}$ of Theorem 10.1.2 on p.213 of [15]).
@@
5 The coincidence of (IR2) and (IR3) in Theorem 3.3 of [2] and (3.20)-(3.22) in Theorem 3.5 of [3]
Now in the studies of the next two propositions, we prove that (IR2) and (IR3) of [2] coincide completely with (3.20)-(3.22) of Theorem 3.5 in [3]. In these proofs we use the notations in [3] and [15].
Proposition 3. The assertion (IR2”) and then (IR2) in [2] coincides precisely with (3.20)-(3.21) in [3].
Proof. Note first by Corollary 6.1.12 and Definition 6.1.13 on p.144 of [15], it yields immediately (4.3)-(4.5). Therefore (IR2”) and then (IR2) in [2] coincides precisely with (3.20)-(3.21) of [3].
@@
Proposition 4. The assertion (IR3”) and then (IR3) in [2] coincides precisely with (3.20)-(3.22) in [3].
Proof. In order to use uniformly the notations in [3] in the following proof, we have changed the notations ${\bf R}^{2m}$, ${\rm Sp}(2m)$, $\ell$ and $\Phi$ in (IR3”) of [2] to ${\bf R}^{2n}$, ${\rm Sp}(2n)$, $m$ and $\gamma$} as those in (3.20) and (3.22) of [3] respectively,
Note first that by the requirement of Theorem 3.3 of [2] on the choice of $k$, the degree $D$ of ${\gamma}_{k}$ is a factor of $k$, and by the requirement (3.30) on $m_{k}$ of Theorem 3.5 of [3], there holds $\frac{m_{k}{\theta}}{\pi}\in{\bf Z}$ whenever $\frac{{\theta}}{\pi}\in{\bf Q}\cap(0,2)$ and $e^{\sqrt{-1}{\theta}}\in{\sigma}(M_{k})$. These two requirements are precisely the same.
By (IR3”) in [2] and (3.22) in [3], to prove Proposition 4, it suffices to prove the following identity for every symplectic path in $C([0,1],{\rm Sp}(2n))$,
@@
where $M_{k}={\gamma}_{k}(1)$, and $Q_{k}(m)$ is given by (3.25) of [3].
@@
where the last equality follows from (3.20) of [3].
@@
where to get the last equality we have applied the fact $S_{M_{k}}^{+}(1)=S_{M_{k}}^{-}(1)$ from Lemma 9.1.6 on p.192 of [15].
Then by the Bott-type formula for splitting numbers which was first proved in [9] by Liu, Long and Zhu (cf. Corollary 9.2.4 on p.201 in [15]), we obtain